OpenStudy (anonymous):

what is the solution to √1 - 4x = 3?

OpenStudy (solomonzelman):

\[\sqrt{1-4x}=3\]

OpenStudy (solomonzelman):

\[\sqrt{1-4x}\color{red} {^2}=3\color{red} {^2}\]\[1-4x=9\]\[1-4x\color{red} {-1}=9\color{red} {-1}\]\[-4x=8\]\[-4x\color{red} {\div (-4)}=8\color{red} {\div (-4)}\]\[x=-2\]

OpenStudy (solomonzelman):

MAKES SENSE?

OpenStudy (anonymous):

thank you, I love how you explains it to me, I have one more.